OpenStudy (anonymous):

which equatio represents the line passing through the point (4-16) and is perpendicular to the line

OpenStudy (anonymous):

use slope point form

OpenStudy (anonymous):

ans do u know that?

OpenStudy (anonymous):

y-y1=m(x-x1) since perpendicular therefore m= -1/m

OpenStudy (anonymous):

\[y-y _{1}=-1/m (x-x _{1}) \]

OpenStudy (anonymous):